2/3 of a number is 3 times more than 1/3 of a number.find the number?
1
Expert's answer
2012-02-14T08:56:48-0500
let's denote the unknown number by x. Then we can construct the following equation:
2/3*x=3*1/3*x
whence x(2/3-1)=0
2/3-1<>0 and therefore the only possible number to satisfy the equation is x=0 and it's the answer to the question.